OpenStudy (anonymous):

factor 9b^2+12bt+4t^2

OpenStudy (anonymous):

Consider:\[(x+y)^2=x^2+2xy+y^2\]If we let:\[x=3b,\ y=2t\]

OpenStudy (anonymous):

Then there is this pattern: \[(3b)^2+2(3b)(2t)+(2t)^2\]

OpenStudy (anonymous):

@gera77 Get it?

OpenStudy (anonymous):

sorry the answer is (3b+2t)(3b+2t) is a special factoring

OpenStudy (anonymous):

\[(3b+2t)(3b+2t) = (3b+2t)^2\]

OpenStudy (anonymous):

that is the answer thank you though

OpenStudy (anonymous):

Ummm, you don't understand what I said?

OpenStudy (anonymous):

yep your answer is right. thanks a lot buddy